If the diameter is 4 inches, and the height is 14 inches, what is the surface area of the cylinder
Answer by bluemockingjay7(41)

formula of the surface area of the cylinder: 2(3.14)rh
diameter = 4/2 = 2 = radius
height = 14
Solution:
surface area=2(3.14)(2)(14)
surface area=2(3.14)(28)
surface area=2(87.92)
surface area=175.84 square inches